OpenStudy (anonymous):

log9 (1/27) -1.5????

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

\[\log_{9} (1/27) = \frac{ \log(1/27 }{ \log(9 }=\frac{ -3 }{ 2 }\]

OpenStudy (anonymous):

thank you (: